OpenStudy (anonymous):

How long will it take to double an amount invested at 8% compunded quarterly?

OpenStudy (anonymous):

I mean 9% .. sorry.

OpenStudy (anonymous):

A=P(1+0.08/4)^4

OpenStudy (anonymous):

That's the answer???

OpenStudy (anonymous):

you have a number for plug in?

OpenStudy (anonymous):

P for principlal money you deposit

OpenStudy (anonymous):

I'm trying to find P??? I don't have any more numbers..

OpenStudy (anonymous):

sorry you said 9%

OpenStudy (anonymous):

if you don have number for P think teacher only need the equation, not the answer number

OpenStudy (anonymous):

A=P(1+0.09/4)^4 that all

OpenStudy (anonymous):

if you have A (total ) you can find pricipal amout, or you have principal amout ,you can find the total

OpenStudy (anonymous):

So it's 1.009030421?

OpenStudy (anonymous):

how you get this number 1.009030421?

OpenStudy (anonymous):

From (1+0.09/4)^4

OpenStudy (anonymous):

you should check it you missing one more number?

OpenStudy (anonymous):

I am not. That's the given problem.

OpenStudy (anonymous):

I missing you said double 2A =P (1+0.09/4)^4 but still not enough information,you may repost question again to see some one have differen answer